# Cosmological simulations of the generation of cluster-scale radio emission from turbulent re-acceleration

### v1

@@ 3.5 Model variations
Finally, we varied the epoch for the first injection of the seed electron population associated with our tracers. This has a negligible effect, unless in the unrealistic case in which all electrons are injected just in the latest 2 Gyr of the simulation, i.e. after the last major merger.
@@ 4 Discussion and conclusions
In conclusion, our study supports the idea that diffuse radio emission in clusters of galaxies can be produced by the turbulent re-acceleration of relativistic electrons. In particular, both ClassicalRH and MegaRH form following major mergers and the resulting dissipation of solenoidal kinetic energy. MegaRH result from a more anisotropic diffuse radio component that is elongated approximately along the direction of the last major merger. They also result from more solenoidal and more supersonic turbulence.
@@ Acknowledgements
FV acknowledges the financial support from the Cariplo ”BREAKTHRU” funds Rif: 2022-2088 CUP J33C22004310003. FV acknowledges the usage of computing at the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for supporting this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC), under projects ”BREAKTHRU”. FV acknowledges the usage of computing time on the LEONARDO HPC System at CINECA, under project ”IsB28_RADGALEO”, with F.V. as Principal Investigator. MB acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and project number 443220636 (DFG research unit FOR 5195: ”Relativistic Jets in Active Galaxies”).

### v2

@@ 3.5 Model variations
Finally, we varied the epoch for the first injection of the seed electron population associated with our tracers. In detail, we resimulated the evolution of radio emissions for tracers, considering injection just before or after the assembly of most of the cluster mass (which is roughly located at a cosmic time $\sim 8.6\rm\leavevmode\nobreak\ Gyr$ since), as well as just before or after the last major merger experienced by the cluster (cosmic time $\sim 12\rm\leavevmode\nobreak\ Gyr$). These variations have overall a negligible effect, unless in the unrealistic case in which all electrons are injected just in the latest 2 Gyr of the simulation, i.e. after the last major merger.
@@ 4 Discussion and conclusions
Another important aspect which can only be explored with a larger sample of simulated clusters, is the role played by the merger mass ratio. While it is at most 1:5 in the cluster studied in this first pilot study, smaller mass ratio can lead to an increased relative amount of turbulent injection. For example, Cassano & Brunetti (2005) estimated with a semi-analytical treatment of ram-pressure stripping that a 1:2 merger induces about twice as a much turbulent energy (relative to the thermal energy), than a 1:5 merger, which would implies a more efficient acceleration of electrons. However, it is difficult to predict also the associated effect of shock dissipation and shock injection of electrons on the formation of diffuse radio emission, and direct and time-dependent simulations are best suited to study such non-linear process in detail, especially in the little explored regime of MegaRH. In conclusion, our study supports the idea that diffuse radio emission in clusters of galaxies can be produced by the turbulent re-acceleration of relativistic electrons. In particular, both ClassicalRH and MegaRH form following major mergers and the resulting dissipation of solenoidal kinetic energy. MegaRH result from a more anisotropic diffuse radio component that is elongated approximately along the direction of the last major merger. They also result from more solenoidal and more supersonic turbulence.
@@ Acknowledgements
We thank our anonymous reviewer for the precious suggestions and comments to the first version of this paper. F.V. has been supported by Fondazione Cariplo and Fondazione CDP, through grant n. Rif: 2022-2088 CUP J33C22004310003 for ”BREAKTHRU” project. F.V. also acknowledges the usage of computing at the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for supporting this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC), under projects ”BREAKTHRU”. FV acknowledges the usage of computing time on the LEONARDO HPC System at CINECA, under project ”IsB28_RADGALEO”, with F.V. as Principal Investigator. MB acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and project number 443220636 (DFG research unit FOR 5195: ”Relativistic Jets in Active Galaxies”).
Appendix A Fokker-Planck method and loss terms
We closely followed the method outlined by Donnert et al. (2013), which we incorporated in the existing version of the ROGER code (Vazza et al. 2023b), to solve the evolution of the number density of relativistic electrons in momentum space, $N(p)$, using the following Fokker-Planck equation:
$\frac{\partial N(p)}{\partial t}$ $=\frac{\partial}{\partial p}\left[H(p)N(p)+D_{\mathrm{pp}}\frac{\partial N(p)}{\partial p}\right]$
$+Q_{\mathrm{inj}}(p,t)$ (3)
where $Q_{\mathrm{inj}}(p,t)$ is a time-dependent source term for the injection of electrons (in our case, only allowed for DSA) and $H(p)$ is the generalised cooling function $H(p)$:
$H(p)$ $=-\frac{2}{p}D_{\mathrm{pp}}+\sum_{i}\frac{dp}{dt}_{i},$ (4)
which combines all relevant losses for the particle population: ionisation and Coulomb losses, adiabatic losses, synchrotron and Inverse Compton losses, as well as the gain by turbulent acceleration, via the $D_{\mathrm{pp}}$ term.
While the acceleration terms have been discussed in the main paper (Sec. 2.2), the loss terms considered in our work are:
  • Coulomb losses, expressed as:
    $\left(\frac{dp}{dt}\right)_{\mathrm{c}}=-3.3\cdot 10^{-29}n_{\mathrm{th}}\left[1+\frac{ln(\gamma/n_{\mathrm{th}})}{75}\right]\ ,$ (5)
    where $n_{\mathrm{th}}$ is the thermal proton density in the medium in $\rm cm^{-3}$ and $\gamma$ is the Lorentz factor.
  • Synchrotron and Inverse Compton losses, expressed as:
    $\left(\frac{dp}{dt}\right)_{\mathrm{rad}}=-4.8\cdot 10^{-4}p^{2}\left[\left(\frac{B_{\mathrm{\mu G}}}{3.2}\right)^{2}+(1+z)^{4}\right]\ ,$ (6)
    where, $z$ is the redshift, $B_{\mathrm{\mu G}}$ is the ICM magnetic field in $\mu\mathrm{G}$ and where we assumed isotropic distribution of momenta with respect to the magnetic field. The term between the square brackets represents the magnetic field strength of the ICM, while the second is the equivalent magnetic field strength that an electron would experience by interacting with the background of CMB photons during an IC scattering process.
  • reversible adiabatic changes (either losses or gains) are modelled through the divergence of the velocity field, which is measured on board of tracers:
    $\left(\frac{dp}{dt}\right)_{\mathrm{ad}}=\frac{2\leavevmode\nobreak\ p}{3}\cdot(\@vec{\nabla}\cdot\@vec{v})\ ,$ (7)
We neglect bremsstrahlung losses since their timescale is significantly larger than the ones of all other loss channels for the ICM physical condition considered here.
We solved the equations based on the finite difference scheme by Chang & Cooper (1970) and with the method by Donnert et al. (2013), which incorporates both systematic and stochastic effects of Fermi II acceleration, together with all most important cooling effects. This scheme ensures particle number conservation by construction, it is suitable for logarithmic grids and it is unconditionally stable, allowing to us to use a small number of momentum bins.